# Quantum Coherent Transport of 1D ballistic states in second order topological insulator Bi4Br4

### v1

@@ V Negative 4-wire resistances
While the behavior of samples contacted from above and from below are similar on many aspects, there are a few clear differences. The first is the presence of a higher interface resistance for bottom contacts, and the second is the absence of a 4-wire negative differential resistance in top-contacted samples.
@@
Indeed, we observe 4-wire negative resistance fluctuations that extend between -6 k and 10 k with a 2-wire resistance of 20-35 k of (Fig. 6). This corresponds to a fraction of the theoretical limit of , predicted for ballistic conduction measured with moderately invasive voltage probes [53]. Indeed, the measured 2-wire resistance of the voltage probe is high, indicating a weakly perturbative probes of the edge states that can favor the appearance of these negative fluctuations. Moreover, contacts from below may allow for the hinge states to be preserved to a certain extent, if they are not cut by a large Pd/Bi alloy region. We emphasize that this prediction of a 4-wire negative differential resistance for ballistic systems has not been reported by many studies. Experiments on single-walled carbon nanotubes have found negative fluctuations reaching the theoretical limit [54], as well as measurements on high-mobility GaAs 2DEGs, which were observed only in a specific field range [55, 56]. Both of these experiments featured exceptionally clean systems with non-invasive voltage probes, which contrasts with the disorder induced at the contacts with Bi4Br4. Therefore, this observation of negative differential resistance in our samples could be a consequence of the expected topological protection of hinge states to elastic disorder.
VI Aharonov-Bohm interferences
@@
In particular the longitudinal magnetoresistance of sample shows a very clear periodicity of around 0.24 T at low temperatures leading to a FFT peak at 4 T-1 (Fig.7a and c). We interpret this as Aharonov-Bohm (AB) interferences [57] between two 1D states of length 6 m and separed by 2.7 nanometers, allowed by the partially coherent contact region, as confirmed by numerical simulations (Fig. 8b and [65]). One possible reason for this particularly clear signature of AB interference in this sample is the small separation between the lateral electrodes, which selects only a few channels that can carry current coherently across the flake without being interrupted by a contact (Fig.7b). We find an exponential decay with temperature of the AB amplitude (Inset of Fig. 7c). This temperature behaviour has been reported in quasi-1D ballistic rings [64], 3D topological insulators [60, 61, 62] and hinge state loops in thin MoTe2 flakes [63]. This dependence is however at odds with the result of Hossain et al. [39] on thin flakes of Bi4Br4, which instead exhibits a power law dependence that is attributed to the extremely long, in effect temperature-independent, phase coherence length of edge states. The functional difference between the decay of the UCFs and AB oscillations with temperature in our samples suggests that their coherence are limited by distinct mechanisms. One would expect the amplitude of AB oscillations to be proportional to the probability of an electron trajectory to complete a loop without phase-breaking event , with the perimeter of the closed trajectory [57]. The exponential decay of the peak amplitude in the FFT with temperature therefore implies a 1/T dependence of the phase coherence length and, taking m, m at 1 K. The fact that we can measure long phase coherence length implies that phase-breaking occurs predominantly within the hinge states, so that the electron trajectories contributing the most to the AB effect do not explore the diffusive regions much, thanks to the very small separation between hinge states (2.7 nm). However this result should be taken with caution since the AB oscillations are measured on the conductance of the combined system of two partially phase coherent disordered contacts (with 100 nm at low temperatures) and highly coherent edge states, whose interference effects are not entirely disentangled. Nevertheless our results emphasize the coherence of electronic transport in relatively long Bi4Br4 flakes.
@@
Strikingly, for device , several of the FFT peaks correspond to a field scale proportional to the length of the considered segment. Hence if the magnetic frequency is rescaled by the segment length (Fig. 9b), it is possible to identify some peaks in the FFT of different segments that correspond to the same rescaled magnetic frequency. We interpret this as AB interferences around surfaces of fixed widths across each segment. We postulate that these interferences originate from coherent 1D hinge states that extend over the entire length of the flake. However they are divided into several segments by the fabricated contacts. Consequently, from segment to segment, there is a constant lateral separation between the hinge states, as illustrated in figure 9a. As the AB effect is visible for all segments, including the longest ones, the phase coherence length of the edge states must be of the order of 5 m or more. However, the limited phase coherence in the contacts region induces an upper cutoff to the lateral distance between 1D channels allowing for AB interferences, such that we do not observe peaks in the Fourier spectrum corresponding to distances greater than a few tens of nanometers. We attribute the absence of a clear magnetic field periodicity in the magneto-conductance signal to the multiplicity of frequencies present in the Fourier transform, which are due to the many 1D coherent channels of the flake that may interfere with each other. Moreover, the magnetic frequencies identified in the FFT spectrum can be related to some specific location of edge states, which are consistent with the crystalline structure of Bi4Br4 [65]. The temperature dependence of the identified peaks exhibits an exponential decay consistent with a 1/T dependence of phase coherence length and a value at 1K similar to the one of sample .
@@ VII Conclusion
We have probed the electronic transport in micrometer-size Bi4Br4 flakes and have demonstrated a strongly anisotropic transport. We have highlighted two distinct types of contributions to low-temperature coherent transport and argue that they suggest of long coherent 1D hinge states contacted to reservoirs via a small coherent disordered region. The first type of contributions yields universal conductance fluctuations and weak-antilocalization and arises from phase coherence in the region close to the contacts, that modulates transmission through the 1D states. The second type of contributions is the periodic Aharonov-Bohm conductance oscillations due to interference between adjacent 1D coherent channels. We emphasize that these effects are due to the very nature of the contacts, which exhibit some phase coherence despite being rather invasive. Both effects indicate that 1D states are ballistic and coherent over several micrometers at low temperatures. The signature of the 1D ballistic states, located at step edges distributed at the surface of the Bi4Br4 crystal, and their interference are thus paradoxically revealed thanks to the coherent diffusive nature of the contact regions. These results argue in favor of crystal-preserving contact methods to achieve the conductance quantization regime in small samples with a reduced number of edge states. The signature of the helical nature of these 1D states in multilayer samples remains to be demonstrated. Furthermore, the use of superconducting contacts to enter the regime of proximity-induced superconductivity is of great interest, with expected perfect Andreev reflection, and would further enhance the potential of Bi4Br4 as a second-order topological insulator.
@@ Acknowledgements.
The authors acknowledge B. Wieder, E. Gerber, T. Wakamura, M. Kociak, M. Aprili, J. Esteve for fruitful discussions, L. Bugaud’s involvement in the initial steps of the project and technical help from S. Autier-Laurent and R. Weil. This work was funded by JSPS KAKENHI (Grant Numbers 21H04652, 21K18181,and 21H05236), JST CREST (Grant Number JPMJCR20B4), the French Renatech network and the European Union through the BALLISTOP ERC 66566 advanced grant.References
- [1] M. Z. Hasan and C. L. Kane, Colloquium: Topological insulators, Rev. Mod. Phys. 82, 3045 (2010).
- [2] X.-L. Qi and S.-C. Zhang, Topological insulators and superconductors, Rev. Mod. Phys. 83, (2011).
- [3] Y. Ando, Topological Insulator Materials, J. Phys. Soc. Jpn. 82, 102001 (2013).
- [4] C. L. Kane and E. J. Mele, Topological Order and the Quantum Spin Hall Effect, Phys. Rev. Lett. 95, 146802 (2005).
- [5] B. A. Bernevig and S.-C. Zhang, Quantum Spin Hall Effect, Phys. Rev. Lett. 96, 106802 (2006).
- [6] F. Schindler, A. M. Cook, M. G. Vergniory, Z. Wang, S. S. P. Parkin, B. A. Bernevig, and T. Neupert, Higher-order topological insulators, Science Advances 4, eaat0346 (2018).
- [7] B. Xie, H.-X. Wang, X. Zhang, P. Zhan, J.-H. Jiang, M. Lu, and Y. Chen, Higher-order band topology, Nat Rev Phys 3, 520 (2021).
- [8] M. König, S. Wiedmann, C. Brüne, A. Roth, H. Buhmann, L. W. Molenkamp, X.-L. Qi, and S.-C. Zhang, Quantum Spin Hall Insulator State in HgTe Quantum Wells, Science 318, 766 (2007).
- [9] Z. Fei, T. Palomaki, S. Wu, W. Zhao, X. Cai, B. Sun, P. Nguyen, J. Finney, X. Xu, and D. H. Cobden, Edge conduction in monolayer WTe2, Nature Phys 13, 7 (2017).
- [10] S. Wu, V. Fatemi, Q. D. Gibson, K. Watanabe, T. Taniguchi, R. J. Cava, and P. Jarillo-Herrero, Observation of the quantum spin Hall effect up to 100 kelvin in a monolayer crystal, Science 359, 76 (2018).
- [11] A. Murani et al., Ballistic edge states in Bismuth nanowires revealed by SQUID interferometry, Nature Communications 8, ncomms15941 (2017).
- [12] F. Schindler et al., Higher-order topology in bismuth, Nature Phys 14, 918 (2018).
- [13] J. Tang et al., Dual quantum spin Hall insulator by density-tuned correlations in TaIrTe4, Nature 628, 515 (2024).
- [14] B. Weber et al., 2024 roadmap on 2D topological insulators, J. Phys. Mater. 7, 022501 (2024).
- [15] J.-J. Zhou, W. Feng, C.-C. Liu, S. Guan, and Y. Yao, Large-Gap Quantum Spin Hall Insulator in Single Layer Bismuth Monobromide Bi 4 Br 4, Nano Lett. 14, 4767 (2014).
- [16] J.-J. Zhou, W. Feng, G.-B. Liu, and Y. Yao, Topological edge states in single- and multi-layer Bi4Br4, New J. Phys. 17, 015004 (2015).
- [17] C.-H. Hsu, X. Zhou, Q. Ma, N. Gedik, A. Bansil, V. M. Pereira, H. Lin, L. Fu, S.-Y. Xu, and T.-R. Chang, Purely rotational symmetry-protected topological crystalline insulator -Bi4Br4, 2D Mater. 6, 031004 (2019).
- [18] F. Tang, H. C. Po, A. Vishwanath, and X. Wan, Efficient topological materials discovery using symmetry indicators, Nat. Phys. 15, 470 (2019).
- [19] C. Yoon, C.-C. Liu, H. Min, and F. Zhang, Quasi-One-Dimensional Higher-Order Topological Insulators, arXiv:2005.14710 [Cond-Mat] (2020).
- [20] K.-S. Lin et al., Spin-resolved topology and partial axion angles in three-dimensional insulators, Nat Commun 15, 550 (2024).
- [21] J. Han, W. Xiao, and Y. Yao, Quasi-one-dimensional topological material Bi4X4(X=Br,I), Advances in Physics: X 7, 2057234 (2022).
- [22] R. Noguchi et al., Evidence for a higher-order topological insulator in a three-dimensional material built from van der Waals stacking of bismuth-halide chains, Nature Materials 1 (2021).
- [23] Yang, M. et al. Large-Gap Quantum Spin Hall State and Temperature-Induced Lifshitz Transition in Bi 4 Br 4. ACS Nano 16, 3036–3044 (2022).
- [24] Zhong, J. et al. Coalescence of multiple topological orders in quasi-one-dimensional bismuth halide chains. Nat Commun 16, 1163 (2025).
- [25] W. Zhao et al., Topological electronic structure and spin texture of quasi-one-dimensional higher-order topological insulator Bi4Br4, Nat Commun 14, 8089 (2023).
- [26] J. Han et al., Optical bulk-boundary dichotomy in a quantum spin Hall insulator, Science Bulletin 68, 417 (2023).
- [27] N. Shumiya et al., Evidence of a room-temperature quantum spin Hall edge state in a higher-order topological insulator, Nat. Mater. 1 (2022).
- [28] J. K. Hofmann et al., Shear-Resistant Topology in Quasi One-Dimensional van Der Waals Material Bi4Br4, arXiv:2411.13320.
- [29] Peng, X. et al. Observation of Topological Edge States on Bi4Br4 Nanowires Grown on TiSe 2 Substrates. J. Phys. Chem. Lett. 12, 10465–10471 (2021).
- [30] D.-Y. Chen, D. Ma, J. Duan, D. Chen, H. Liu, J. Han, and Y. Yao, Quantum transport evidence of boundary states and Lifshitz transition in , Phys. Rev. B 106, 075206 (2022).
- [31] H. von Benda, A. Simon, and W. Bauhofer, Zur Kenntnis von BiBr und BiBr1,167, Zeitschrift für anorganische und allgemeine Chemie 438, 53 (1978).
- [32] T. G. Filatova, P. V. Gurin, L. Kloo, V. A. Kulbachinskii, A. N. Kuznetsov, V. G. Kytin, M. Lindsjo, and B. A. Popovkin, Electronic structure, galvanomagnetic and magnetic properties of the bismuth subhalides Bi4I4 and Bi4Br4, Journal of Solid State Chemistry 180, 1103 (2007).
- [33] X. Li et al., Pressure-induced phase transitions and superconductivity in a quasi–1-dimensional topological crystalline insulator -Bi4Br4, PNAS 116, 17696 (2019).
- [34] Z. Gong et al., Br-Vacancies Induced Variable Ranging Hopping Conduction in High-Order Topological Insulator Bi4Br4, Small Methods, 2400517 (2024).
- [35] J. Zhong, M. Yang, F. Ye, C. Liu, J. Wang, J. Wang, W. Hao, J. Zhuang, and Y. Du, Facet-dependent Electronic Quantum Diffusion in the High-Order Topological Insulator , Phys. Rev. Applied 17, 064017 (2022).
- [36] Zhong, J. et al. Observation of Anomalous Planar Hall Effect Induced by One-Dimensional Weak Antilocalization. ACS Nano acsnano.3c10120 (2024)
- [37] L. Qiao et al., Ultralong Single-Crystal -Bi4Br4 Nanobelts with a High Current Carrying Capacity by Mechanical Exfoliation, J. Phys. Chem. C 125, 22312 (2021).
- [38] S.-L. Wu, Z.-H. Ren, Y.-Q. Zhang, Y.-K. Li, J.-F. Han, J.-X. Duan, Z.-W. Wang, C.-Z. Li, and Y.-G. Yao, Gate-tunable transport in van der Waals topological insulator Bi4Br4 nanobelts, J. Phys.: Condens. Matter 35, 234001 (2023).
- [39] M. S. Hossain et al., Quantum transport response of topological hinge modes, Nat. Phys. 1 (2024).
- [40] R. Noguchi et al., A weak topological insulator state in quasi-one-dimensional bismuth iodide, Nature 566, 7745 (2019).
- [41] Bapat, R. B, Resistance matrix of a weighted graph, Communications in Mathematical and in Computer Chemistry 50, 73‑82 (2004).
- [42] Devriendt Karel, Graph Geometry from Effective Resistances, PhD Thesis, Trinity College, 2022.
- [43] Datta, S. Electronic Transport in Mesoscopic Systems. (Cambridge Univ. Press, Cambridge, 2009).
- [44] E. Akkermans and G. Montambaux, Mesoscopic Physics of Electrons and Photons, Cambridge University Press (2007).
- [45] S. Hikami, A. I. Larkin, and Y. Nagaoka, Spin-Orbit Interaction and Magnetoresistance in the Two Dimensional Random System, Progress of Theoretical Physics 63, 707 (1980).
- [46] H. Steinberg, J.-B. Laloë, V. Fatemi, J. S. Moodera, and P. Jarillo-Herrero, Electrically tunable surface-to-bulk coherent coupling in topological insulator thin films, Phys. Rev. B 84, 233101 (2011).
- [47] J. Chen, X. Y. He, K. H. Wu, Z. Q. Ji, L. Lu, J. R. Shi, J. H. Smet, and Y. Q. Li, Tunable surface conductivity in Bi 2 Se 3 revealed in diffusive electron transport, Phys. Rev. B 83, 241304 (2011).
- [48] P. A. Lee, A. D. Stone, and H. Fukuyama, Universal conductance fluctuations in metals: Effects of finite temperature, interactions, and magnetic field, Phys. Rev. B 35, 1039 (1987).
- [49] P. Mohanty and R. A. Webb, High-Field Measurements of Electron Decoherence Time in Metallic Nanowires: Switching off Magnetic Impurity Spins, Phys. Rev. Lett. 91, 066604 (2003).
- [50] C. W. J. Beenakker and H. Van Houten, Flux-cancellation effect on narrow-channel magnetoresistance fluctuations, Phys. Rev. B 37, 6544 (1988).
- [51] Groth, C. W., Wimmer, M., Akhmerov, A. R. and Waintal, X. Kwant: a software package for quantum transport. New J. Phys. 16, 063065 (2014).
- [52] M. Büttiker, Absence of backscattering in the quantum Hall effect in multiprobe conductors, Phys. Rev. B 38, 9375 (1988).
- [53] M. Büttiker, Symmetry of electrical conduction, IBM J. Res. Dev. 32, 317 (1988).
- [54] B. Gao, Y. F. Chen, M. S. Fuhrer, D. C. Glattli, and A. Bachtold, Four-Point Resistance of Individual Single-Wall Carbon Nanotubes, Phys. Rev. Lett. 95, 196802 (2005).
- [55] G. Timp, A. M. Chang, P. Mankiewich, R. Behringer, J. E. Cunningham, T. Y. Chang, and R. E. Howard, Quantum transport in an electron-wave guide, Phys. Rev. Lett. 59, 732 (1987).
- [56] de Picciotto, R., Stormer, H. L., Pfeiffer, L. N., Baldwin, K. W. and West, K. W. Four-terminal resistance of a ballistic quantum wire. Nature 411, 51–54 (2001).
- [57] S. Washburn and R. A. Webb, Aharonov-Bohm effect in normal metal quantum coherence and transport, Advances in Physics 35, 375 (1986).
- [58] Altshuler, B. L., Aronov, A. G. and Spivak, B. Z. The Aharonov–Bohm effect in disordered conductors. JETP Lett. 33, 94–97 (1981).
- [59] Washburn, S., Umbach, C. P., Laibowitz, R. B. and Webb, R. A. Temperature dependence of the normal-metal Aharonov–Bohm effect. Phys. Rev. B 32, 4789–4792 (1985).
- [60] Dufouleur, J. et al. Quasiballistic Transport of Dirac Fermions in a Bi 2 Se 3 Nanowire. Phys. Rev. Lett. 110, 186806 (2013).
- [61] Jauregui, L. A., Pettes, M. T., Rokhinson, L. P., Shi, L. and Chen, Y. P. Magnetic field-induced helical mode and topological transitions in a topological insulator nanoribbon. Nature Nanotech 11, 345–351 (2016).
- [62] G. Behner et al., Aharonov-Bohm Interference and Phase-Coherent Surface-State Transport in Topological Insulator Rings, Nano Lett. 23, 6347 (2023).
- [63] Pan, Z.-C. et al. Altshuler-Aronov-Spivak interference of one-dimensional helical edge states in . Phys. Rev. B 107, 045411 (2023).
- [64] A. E. Hansen, A. Kristensen, S. Pedersen, C. B. Sørensen, and P. E. Lindelof, Mesoscopic decoherence in Aharonov-Bohm rings, Phys. Rev. B 64, 045327 (2001).
- [65] See Supplemental Material for a discussion of the crystalline quality of Bi4Br4 flakes, resistance networks of two other samples at different temperatures, the length dependence of resistance for samples and , we expand on the temperature dependence of the Aharonov-Bohm effect for sample , details of the numerical simulations presented in the main text, the fitting of the magnetoconductance using the formula for 1D weak anti-localization, and finally a presentation of zero-bias anomaly data and comparison with theoretical predictions.

### v2

@@ V Negative four-wire resistance
While the behavior of samples contacted from above and from below are similar on many aspects, there are a few clear differences. The first is the presence of a higher interface resistance for bottom contacts, and the second is the absence of a four-wire negative differential resistance in top-contacted samples.
@@
Indeed, we observe negative four-wire resistance fluctuations that extend between -6 and 10 k with a two-wire resistance of 20-35 k for (Fig. 6). This corresponds to a sizable fraction of the theoretical limit of , predicted for ballistic conduction measured with moderately invasive voltage probes [76]. Indeed, the two-wire resistance measured through the voltage probes indicated in blue in Fig. 6 is high (approximately 150 k), which may indicate a weakly perturbative contact to the edge states that can favor the appearance of these negative fluctuations. Indeed, contacts from below may allow for the hinge states to be preserved to a certain extent, if they are not cut by a large Pd/Bi alloy region (Fig. 1). We emphasize that this prediction of a four-wire negative differential resistance for ballistic systems has not been reported by many studies. Experiments on single-walled carbon nanotubes have found negative fluctuations reaching the theoretical limit [77], as well as measurements on high-mobility GaAs 2DEGs, which were observed only in a specific field range [78, 79]. Both of these experiments featured exceptionally clean systems with noninvasive voltage probes, which contrasts with the disorder induced at the contacts with Bi4Br4. Therefore, this observation of negative differential resistance in our samples could be a consequence of the expected topological protection of hinge states against elastic disorder.
VI Aharonov-Bohm interference
@@
In particular the longitudinal magnetoresistance of sample shows a very clear periodicity of around 0.24 T at low temperatures leading to a FFT peak at 4 T-1 [Fig.7(a) and 7(c)]. Note that the periodicity of the oscillations does not depend on temperature but the amplitude does (Sec. 4 of Supplemental Material [41]). We interpret this as Aharonov-Bohm (AB) interference [80] between two 1D states of length 6 m and separated by 2.7 nm, allowed by the partially coherent contact region, as confirmed by numerical simulations [Fig. 8(b) and Sec. 5 of Supplemental Material [41]). One possible reason for this particularly clear signature of AB interference in this sample is the small separation between the lateral electrodes, which selects only a few channels that can carry current coherently across the flake without being interrupted by a contact [Fig.7(b)]. The good visibility of the AB oscillations is also favored by the small lateral distance between edge states, comparable to the mean free path in the diffusive contact [Sec. 5 of Supplemental Material [41]]. Besides, we find an exponential decay with temperature of the AB amplitude [inset in Fig. 7(c)). This temperature behavior has been reported in quasi-1D ballistic rings [85], 3D topological insulators [81, 82, 83] and hinge state loops in thin MoTe2 flakes [84]. This dependence is, however, at odds with the result of Hossain et al. [39] on thin flakes of Bi4Br4, which instead exhibits a power-law dependence that is attributed to the extremely long, effectively temperature-independent, phase coherence length of edge states. The functional difference between the decay of the UCFs and AB oscillations with temperature in our samples suggests that their coherence is limited by distinct mechanisms. One would expect the amplitude of AB oscillations to be proportional to the probability of an electron trajectory to complete a loop without phase-breaking event , with the perimeter of the closed trajectory [80]. The exponential decay of the peak amplitude in the FFT with temperature, therefore, implies a dependence of the phase coherence length and taking m, m at 1 K. The fact that we can measure such a long phase coherence length implies that phase breaking occurs predominantly within the hinge states, so that the electron trajectories contributing the most to the AB effect do not explore the diffusive regions much, thanks to the very small separation between hinge states (2.7 nm). Our results emphasize the coherence of electronic transport in relatively long Bi4Br4 flakes.
@@
Strikingly, for device , several of the FFT peaks correspond to a field scale proportional to the length of the considered segment. Hence, if the magnetic frequency is rescaled by the segment length [Fig. 9(b)], it is possible to identify some peaks in the FFT of different segments that correspond to the same rescaled magnetic frequency. We interpret this as AB interference around surfaces of fixed widths across each segment. We postulate that these oscillations originate from coherent 1D hinge states that extend over the entire length of the flake, which were subsequently divided into several segments by the fabricated contacts. Consequently, from segment to segment, there is a constant lateral separation between the hinge states, as illustrated in Fig. 9(a). In fact, the magnetic frequencies identified in the FFT spectrum can be related to some specific configurations of hinge states that are consistent with the crystalline structure of Bi4Br4 [Fig. 9(b)]. As the AB effect is visible for all segments, including the longest ones, the phase coherence length of the edge states must be of the order of 5 m or more. However, the limited phase coherence in the contact regions induces an upper cutoff to the lateral distance between 1D channels allowing for AB interferences, such that we do not observe peaks in the Fourier spectrum corresponding to distances greater than a few tens of nanometers. Moreover, we attribute the absence of a clear magnetic field periodicity in the magnetoconductance signal to the multiplicity of frequencies present in the Fourier transform, which are due to the many 1D coherent channels of the flake that may interfere with each other. The temperature dependence of the identified peaks exhibits an exponential decay consistent with a dependence of phase coherence length and a value at 1 K similar to the one of sample . Moreover the existence of AB interference in all segments of up to 5 µm is consistent with long phase coherence length.
@@ VII Conclusion
We have probed the electronic transport in micrometer-size Bi4Br4 flakes and have demonstrated a strongly anisotropic transport. We have highlighted two distinct types of contributions to low-temperature coherent transport and argue that they reveal the existence of long coherent 1D hinge states contacted to reservoirs via a small coherent disordered region. The first type of contribution yields universal conductance fluctuations and weak antilocalization, arising from phase coherence in the region close to the contacts, which modulates transmission through the 1D states. The second type of contribution is the periodic Aharonov-Bohm conductance oscillations due to interference between adjacent 1D coherent channels. We emphasize that both of these effects are due to the very nature of the contacts, which exhibit some phase coherence despite being rather invasive. These effects indicate that the 1D states are ballistic and coherent over several micrometers at low temperatures. The signature of the 1D ballistic states, located at step edges distributed at the surface of the Bi4Br4 crystal, and their interference are thus paradoxically revealed thanks to the coherent diffusive nature of the contact regions. It would be interesting to achieve the opposite limit of clean, crystal-preserving contacts in small samples with a reduced number of edge states, in order to access the conductance quantization regime. Also, the signature of the helical nature of these 1D states in multilayer samples remains to be demonstrated. Lastly, the use of superconducting contacts to enter the regime of proximity-induced superconductivity is of great interest, with expected perfect Andreev reflection, and would further enhance the potential of Bi4Br4 as a second-order topological insulator.
@@ Acknowledgements.
The authors acknowledge B. Wieder, E. Gerber, T. Wakamura, M. Kociak, M. Aprili and J. Esteve for fruitful discussions, L. Bugaud’s involvement in the initial steps of the project, and technical help from S. Autier-Laurent and R. Weil. This work was funded by JSPS KAKENHI (Grants No. 21H04652, No. 21K18181 and No. 21H05236), JST CREST (Grant No. JPMJCR20B4), the French Renatech network, the CPER Hauts de France project IMITECH, the Métropole Européenne de Lille, and the European Union through the BALLISTOP ERC 66566 advanced grant.DATA AVAILABILITY
The data that support the findings of this article are openly available [86].